determine the horizontal force P required to start moving the 500N block up the inclined surface.assume static friction is 0.3 and angle is 30


1
Expert's answer
2020-09-21T11:13:25-0400

As per the question

weight of block ,W=500N

Angle on inclination, "\theta=30\degree"

cofficient of friction, "u_s=0.3"


As the block is moving up the inclined surface ,therefore the equation became

"pcos\theta=Wsin\theta+f_s"

"pcos\theta=Wsin\theta+u_sR"

Where R is the Normal reaction to the block from inclined plane

"pcos30=500sin30+0.3(Wcos\theta+psin\theta)"

"0.866p=250+0.3(433+psin30)"

0.866p=250+"0.3\times433+0.3\times p\times Sin30"

"0.866p=250+129.9+0.15p"

"0.866p-0.15p=379.9"

0.716p=379.9

P="\frac{379.9}{0.716}"

=530.58N